Please answer asap! thank you!Rio Bus Tours has incurred the following bus maintenance costs during the recent tourist season. (The real is Brazils nationComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Develop a formula to exprRequired 1 Required 2 Required 3 Predict the level of maintenance cost that would be incurred during a month when 37,000 tour
Answer #1
Highest level of activity=30,800
Cost at highest level of activity = 20,450
Lowest level of activity=12,800
Cost at highest level of activity = 17,300
Variable cost per unit =Difference in cost/Difference in activity level=(20450-17300)/(30800-12800)=0.175
Fixed cost =20450-(30800*0.175)=15060
2
Cost formula = 15,060+0.175X
where X is the kilometres travelled by tour buses
3
Cost at 37000 km =15060+(37000*0.175)=21,535

if any doubt please ,mention in comment